On a canoe trip, Rita paddled upstream (against the current) at an average speed of 2 mi/h relative to the riverbank. On the return trip downstream (with the current), her average speed was 3 mi/h. Find Rita's paddling speed in still water and the speed of the river's current.
Answer by richwmiller(17219)

2+3=5
5/2=2.5
2.5 mph in still water
against the current
3-2=1
1/2=.5
current .5 mph